# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v2

@@ Abstract
Spiking Graph Networks (SGNs) have garnered significant attraction from both researchers and industry due to their ability to address energy consumption challenges in graph classification. However, SGNs are only effective for in-distribution data and cannot tackle out-of-distribution data. In this paper, we first propose the domain adaptation problem in SGNs, and introduce a novel framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA). The proposed DeSGDA addresses the spiking graph domain adaptation problem by three aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation. First, we introduce the personalized spiking representation method for generating degree-dependent spiking signals. Specifically, the threshold of triggering a spike is determined by the node degree, allowing this personalized approach to capture more expressive information for classification. Then, we propose the graph feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator. Such an alignment module can efficiently maintain high performance and low energy consumption in the case of inconsistent distribution. Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance graph classification performance. Extensive experiments on benchmark datasets validate the superiority of the proposed DeSGDA compared with competitive baselines.
@@ 1 Introduction
Spiking Graph Networks (SGNs) Xu et al. (2021b); Zhu et al. (2022) are a specialized type of artificial neural network engineered to process graph information by mimicking the human brain. SGNs transform static and real-valued graph features into discrete spikes by simulating neurons’ charging and discharging cycles, facilitating spike-based representations for graph node classification. Notably, SGNs excel in capturing semantic spiking representations with low energy consumption, which proves advantageous for event-based processing tasks Yao et al. (2021); Rosenfeld et al. (2021); Chen et al. (2023a, b) such as object recognition Gu et al. (2020); Li et al. (2021b); Liu et al. (2023b); Yi et al. (2024), real-time data analysis Bauer et al. (2019); Zhu et al. (2020); Baronig et al. (2024), and graph classification Li et al. (2023); Yao et al. (2023b); Meng et al. (2024); Ju et al. (2024).
@@
Currently, SGNs are usually tested within the same distribution as the training dataset Li et al. (2023); Yin et al. (2024); Duan et al. (2024). However, in realistic scenarios, the testing set can have different distributions from the training set, and such a distribution shift may lead to a degradation in performance. For instance, Electroencephalography (EEG) data Binnie and Prior (1994); Biasiucci et al. (2019); Peng et al. (2022), typically represented as a graph structure with nodes for neurons and edges for connections, is ideally processed by bio-inspired SGNs that mimic neuronal charging and discharging. Despite the suitability, EEGs often exhibit varying distributions over time or among different groups Zhao et al. (2020, 2021); Wang et al. (2022), leading to suboptimal performance of models trained on specific distributions when applied to others. This significant issue underscores the necessity of exploring domain adaptation for spiking graphs. Traditionally, SNNs transfer learning methods Zhan et al. (2021); Zhang et al. (2021); Zhan et al. (2024); Guo et al. (2024) have been applied in event-based or computer vision scenarios. However, there’s no existing research on spiking graph domain adaptation.
@@
Designing an effective spiking graph domain adaptation framework for classification is non-trivial due to the following major challenges: (1) How to meticulously design an SGN under the circumstance of domain shift? SGNs usually utilize a global threshold for the firing of each node Xu et al. (2021a); Hagenaars et al. (2021); Yin et al. (2024); Zhao et al. (2024). However, we observe that the degree of each node influences the difficulty of triggering spikes. Specifically, nodes with high degrees can integrate more information from neighbors, making it easier for membrane potential to accumulate and trigger a spike. Conversely, nodes with lower degrees are more challenging to reach the firing threshold, denoted as the inflexible architecture challenge. (2) How to design a framework that effectively addresses spiking graph domain adaptation for classification? Current research primarily focuses on graph node classification within the same distribution Li et al. (2023); Yao et al. (2023a); Duan et al. (2024); Cai et al. (2024). However, spike-based graph classification under domain shift remains unexplored. (3) How to guarantee the stability of the proposed framework? Though some works have been proposed to address the spiking transfer learning challenges Zhan et al. (2021); Zhang et al. (2021); Liu et al. (2024a); Zhan et al. (2024), there is still no theoretical research on spiking graphs under domain shift.
@@
To tackle these challenges, we propose a framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA), which comprises three components: degree-aware personalized spiking representation, graph feature distribution alignment, and pseudo-label distillation. To address the first challenge, we establish variable node thresholds based on their degrees. By adaptively updating these thresholds, we can achieve a more expressive and personalized spiking representation for each node. Then, we introduce a adversarial feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator. To further enhance performance, we extract consistent predictions from different spaces to generate reliable pseudo-labels. Additionally, to explore the generalization ability of the proposed DeSGDA, we first propose the error bound for spiking graph domain adaptation and demonstrate that our pseudo-label distillation module effectively reduces this upper bound.
@@
Our contributions can be summarized as follows: (1) Problem Formulation: We first introduce the problem of spiking graph domain adaptation for classification, which is non-trivial due to the challenges of the inflexible architecture of SGNs and theoretical deficiency. (2) Novel Architecture: We propose DeSGDA, a framework that efficiently learns personalized spiking representations for nodes using degree-aware thresholds and aligns domain distributions through adversarial training on membrane potential. Furthermore, we utilize pseudo-label distillation to improve the performance further. (3) Theoretical Analysis: To guarantee the stability of DeSGDA, we provide theoretical proof of the error bound for spiking graph domain adaptation. Furthermore, we demonstrate that DeSGDA maintains a lower theoretical bound than standard spiking graph domain adaptation through the effective use of the pseudo-label distillation module. (4) Extensive Experiments. We evaluate the proposed DeSGDA on extensive spiking graph domain adaptation learning datasets, which shows that our proposed DeSGDA outperforms the variety of state-of-the-art methods.
@@ 2 Related work
Spiking Graph Networks (SGNs). SGNs are a specialized type of neural network that combines Spiking Neural Networks (SNNs) with Graph Neural Networks (GNNs), preserving energy efficiency while achieving competitive performance in various graph tasks Li et al. (2023); Yao et al. (2023a); Qiu et al. (2024); Duan et al. (2024). Existing research on SGNs focuses on capturing the dynamic temporal information contained within graphs and enhancing model scalability. For instance, Xu et al. (2021a) utilizes spatial-temporal feature normalization within SNNs to effectively process dynamic graph data, ensuring robust learning and improved predictive performance. Zhao et al. (2024) propose a method that dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real time. Additionally, Yin et al. (2024) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task. However, existing methods still suffer from data distribution shift issues when training and testing data come from different domains, resulting in degraded model performance and generalization. To address this, we propose a novel domain adaptation method based on SGNs to tackle these challenges.
Spiking Transfer Learning. Spiking transfer learning focuses on adjusting SNNs to handle data distribution shifts across various domains effectively. Recent advances in spiking transfer learning have been extensively applied in vision tasks, enhancing model performance while maintaining energy efficiency Zhan et al. (2021); Zhang et al. (2021); Zhan et al. (2024). For instance, Guo et al. (2024) leverages a Jaccard attention mechanism within SNNs to effectively adapt to target domains without requiring source domain data. Similarly, He et al. (2024) facilitates the transfer of learned representations from static to dynamic event-based domains by adapting SNNs to process temporal information. Additionally, Zhan et al. (2024) converts RGB images into spike-based neuromorphic data, enabling SNNs to process visual information across various domains efficiently. However, the difficulty of graph topologies makes it infeasible to apply spiking transfer learning to SGNs directly. To this end, we introduce a specialized domain adaptation method tailored for SGNs.
@@ 3 Preliminaries
(You et al., 2023) Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in (Redko et al., 2017), and the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, with probability at least $1-\delta$ the following inequality holds:
$\epsilon_{T}(h,\hat{h})\leq$ $\hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log(\frac{eN_{S}}{d})+\frac{1}{N_{S}}\log(\frac{1}{\delta})}+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$. The first Wasserstein distance is defined as (Villani et al., 2009): $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$.
@@
Spiking Graph Networks. In contrast to traditional artificial neural networks, SGNs (Xu et al., 2021a; Zhu et al., 2022) convert input data into binary spikes over time, with each neuron in the SGNs maintaining a membrane potential that accumulates input spikes. A spike is produced as an output when the membrane potential exceeds a threshold, which is formulated as:
$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum_{j}w_{ij}\mathcal{A}\left(A,s_{\tau,j}\right)+b,\quad s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),$ (1)
where $\mathbb{H}(x)$ is the Heaviside function, which is the non-differentiable spiking function. $\mathcal{A}$ is the graph aggregation operation, and $A$ is the adjacency matrix of graph. $s_{\tau,i}$ denotes the binary spike train of neuron $i$, and $\lambda$ is the constant. $w_{ij}$ and $b$ are the weights and bias of each neuron.
@@ 4 Methodology
Figure 1: An overview of the proposed DeSGDA. DeSGDA first utilizes different thresholds for different degree nodes to obtain personalized spiking representation. To align domain distribution, we use adversarial training on membrane potential against domain discrimination. Besides, DeSGDA applies the pseudo-label to enhance model performance.
@@ 4.1 Degree-aware Personalized Spiking Representation
Specifically, we first $set$ all the degrees of nodes in the source domain graphs, i.e., $D^{s}=set(D_{1}^{s}\cup\cdots\cup D_{N_{s}}^{s})$, where $D_{i}$ denotes the degree set of graph $G_{i}^{s}$, and $set(\cdot)$ operation is an unordered sequence of non-repeating elements. Considering that low-degree nodes are more challenging to trigger while high-degree nodes trigger more easily, we propose setting higher thresholds for high-degree nodes and lower thresholds for low-degree nodes, which is formulated as:
$s_{\tau}^{d_{i}^{s}}=\mathbb{H}(u_{\tau}-V_{th}^{d_{i}^{s}}),\quad S^{d_{i}^{s}}=avg(s_{\tau}^{d_{i}^{s}}),\quad V_{th}^{d_{i}^{s}}=(1-\alpha)V_{th}^{d_{i}^{s}}+\alpha S^{d_{i}^{s}},$ (2)
where $V_{th}^{d_{i}^{s}}$ is the threshold of degree $d_{i}^{s}\in D^{s}$, initially set to $V_{th}$, and $\alpha$ is a hyper-parameter. The $avg(\cdot)$ operation takes the average of spiking representation with degree $d_{i}^{s}$. Consequently, high-degree nodes tend to achieve high $S^{d_{i}^{s}}$, which leads to an iterative increase in the threshold corresponding to degree $d_{i}^{s}$ and conversely for lower-degree nodes.
With different thresholds for different node degrees, we can obtain the personalized node spiking representation $\mathbf{s}_{v\in G_{i}^{s}}^{d_{j}^{s}}$. Then, we summarize all node representations with a readout function into the graph-level representation and output the prediction with a multi-layer perception (MLP) classifier:
$\mathbf{s}_{i}=\operatorname{READOUT}\left(\left\{\mathbf{s}_{v}^{d_{j}^{s}}\right\}_{v\in G_{i}^{s}}\right),\quad\hat{y}_{i}^{s}=H(\mathbf{s}_{i}),$ (3)
@@
However, in the scenario of domain adaptation, two significant issues remain unresolved. The first issue is that degrees in the target domain may be unseen in the source, rendering the thresholds ineffective for these degrees. The second is that the thresholds in the target may differ from those in the source, simply applying the source domain’s thresholds could lead to performance degradation.
To alleviate the first issue, we initialize the threshold $V_{th}^{d_{i}^{t}}$ with the same value, where $d_{i}^{t}\notin D^{s}$. Then, with the training process of adversarial alignment, we iteratively update the threshold for degree $d_{i}^{t}$ with Eq. 2. To address the second issue, we incorporate the pseudo-label distillation module in Section 4.3 to guide the update of source degree thresholds on the target domain.

### v3

@@ Abstract
Spiking Graph Networks (SGNs) have garnered significant attraction from researchers and industry due to their ability to address energy consumption challenges in graph classification. However, SGNs are only effective for in-distribution data and cannot tackle out-of-distribution data. In this paper, we first propose the domain adaptation problem in SGNs, and introduce a novel framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA). The proposed DeSGDA addresses the spiking graph domain adaptation problem in three aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation. First, we introduce the personalized spiking representation method for generating degree-dependent spiking signals. Specifically, the node degree determines the threshold of triggering a spike, allowing this personalized approach to capture more expressive information for classification. Then, we propose the graph feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator. Such an alignment module can efficiently maintain high performance and low energy consumption in the case of inconsistent distribution. Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance classification performance. Extensive experiments on benchmark datasets validate the superiority of the DeSGDA compared with competitive baselines.
@@ 1 Introduction
Spiking Graph Networks (SGNs) Zhu et al. (2022); Xu et al. (2021b) are a specialized artificial neural network engineered to process graph information by mimicking the human brain. SGNs transform static and real-valued graph features into discrete spikes by simulating neurons’ charging and discharging cycles, facilitating spike-based representations for graph node classification. Notably, SGNs excel in capturing semantic spiking representations with low energy consumption, which proves advantageous for event-based processing tasks Yao et al. (2021); Wang et al. (2024) such as object recognition Gu et al. (2020); Li et al. (2021b); Chen et al. (2024); Yin et al. (2024a), real-time data analysis Zhu et al. (2020); Bauer et al. (2019); Yin et al. (2023b), and graph classification Li et al. (2023); Zhu et al. (2022); Liu et al. (2024c); Yin et al. (2022a).
@@
Currently, SGNs are usually tested within the same distribution as the training dataset Li et al. (2023); Yin et al. (2024a); Duan et al. (2024). However, in realistic scenarios, the testing set can have different distributions from the training set, and such a distribution shift may degrade performance. For instance, Electroencephalography (EEG) data Binnie and Prior (1994); Biasiucci et al. (2019), typically represented as a graph structure with nodes for neurons and edges for connections, is ideally processed by bio-inspired SGNs that mimic neuronal charging and discharging. Despite the suitability, EEGs often exhibit varying distributions over time or among different groups Zhao et al. (2020, 2021); Wang et al. (2022), leading to suboptimal performance of models trained on specific distributions when applied to others. This significant issue underscores the necessity of exploring domain adaptation for spiking graphs. Traditionally, SNNs transfer learning methods Zhan et al. (2021); Yin et al. (2023c, 2024b); Pang et al. (2023) have been applied in event-based or computer vision scenarios. However, there’s no existing research on spiking graph domain adaptation.
@@
In this paper, we address the development of energy-efficient SGNs tailored for scenarios involving distribution shifts. Both domain adaptation and SGNs are particularly well-suited for real-world applications where data distributions vary across environments, and efficient processing of graph-structured, dynamic data under resource constraints is crucial. These challenges are common across numerous fields that require solutions capable of handling distribution shifts while minimizing energy consumption. However, designing an effective spiking graph domain adaptation framework is non-trivial due to the following major challenges: (1) How to meticulously design an SGN under the circumstance of domain shift? SGNs usually utilize a global threshold for the firing of each node Xu et al. (2021a); Yin et al. (2024a); Zhao et al. (2024). However, we observe that the degree of each node influences the difficulty of triggering spikes. Specifically, nodes with high degrees can integrate more information from neighbors, making it easier for membrane potential to accumulate and trigger a spike. Conversely, nodes with lower degrees are more challenging to reach the firing threshold, denoted as the inflexible architecture challenge. (2) How to design a framework that effectively addresses spiking graph domain adaptation for classification? Current research primarily focuses on graph node classification within the same distribution Li et al. (2023); Yao et al. (2023a); Duan et al. (2024). However, spike-based graph classification under domain shift remains unexplored. (3) How to guarantee the stability of the proposed framework? Though some works have been proposed to address the spiking transfer learning challenges Zhan et al. (2021); Zhang et al. (2021); Zhan et al. (2024), there is still no theoretical research on spiking graphs under domain shift.
@@
To tackle these challenges, we propose a framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA), which comprises three components: degree-aware personalized spiking representation, graph feature distribution alignment, and pseudo-label distillation. To address the first challenge, we establish variable node thresholds based on their degrees. By adaptively updating these thresholds, we achieve a more expressive and personalized spiking representation for each node. Then, we introduce an adversarial feature distribution alignment module trained using membrane potential against a domain discriminator. To further enhance performance, we extract consistent predictions from different spaces to generate reliable pseudo-labels. Additionally, to explore the generalization ability of the proposed DeSGDA, we propose the error bound for spiking graph domain adaptation and demonstrate that our pseudo-label distillation module effectively reduces this upper bound. In summary, we utilize simple yet effective techniques to address a novel problem while providing an insightful analysis of the background mechanisms and capabilities of our proposed method.
Our contributions can be summarized as follows:
  • Problem Formulation: We first introduce the problem of spiking graph domain adaptation for classification, which is non-trivial due to the challenges of the inflexible architecture of SGNs and theoretical deficiency.
@@
  • Novel Architecture: We propose DeSGDA, a framework that efficiently learns personalized spiking representations for nodes using degree-aware thresholds and aligns domain distributions through adversarial training on membrane potential. Furthermore, we utilize pseudo-label distillation to improve the performance further.
  • Theoretical Analysis: To guarantee the stability of DeSGDA, we provide theoretical proof of the error bound for spiking graph domain adaptation. Furthermore, we demonstrate that DeSGDA maintains a lower theoretical bound than standard spiking graph domain adaptation through the use of the pseudo-label distillation module.
  • Extensive Experiments. We evaluate the proposed DeSGDA on extensive spiking graph domain adaptation learning datasets, which shows that our proposed DeSGDA outperforms the variety of state-of-the-art methods.
@@ 2 Related work
Figure 1: An overview of the proposed DeSGDA. To achieve personalized spiking representations, DeSGDA employs adaptive thresholds based on node degrees, enabling the generation of tailored spiking representations. To align domain distributions, DeSGDA leverages adversarial training on membrane potentials to counter domain discrimination. Furthermore, DeSGDA utilizes pseudo-labeling to identify and select reliable samples, thereby enhancing overall model performance.
Spiking Graph Networks (SGNs). SGNs are a specialized type of neural network that combines Spiking Neural Networks (SNNs) with Graph Neural Networks (GNNs), preserving energy efficiency while achieving competitive performance in various graph tasks Li et al. (2023); Yao et al. (2023a, b); Duan et al. (2024); Liang et al. (2023). Existing research on SGNs focuses on capturing the dynamic temporal information within graphs and enhancing scalability. For instance, Xu et al. (2021a) utilizes spatial-temporal feature normalization within SNNs to effectively process dynamic graph data, ensuring robust learning and improved predictive performance. Zhao et al. (2024) propose a method that dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real time. Additionally, Yin et al. (2024a) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task. However, existing methods still suffer from data distribution shift issues when training and testing data come from different domains, resulting in degraded performance and generalization. To address this, we propose a novel domain adaptation method based on SGNs to tackle these challenges.
Spiking Transfer Learning. Spiking transfer learning focuses on adjusting SNNs to handle data distribution shifts across various domains more effectively. Recent advances in spiking transfer learning have been extensively applied in vision tasks, significantly enhancing model performance while maintaining energy efficiency Zhan et al. (2021); Zhang et al. (2021); Zhan et al. (2024). For instance, Guo et al. (2024) leverages a Jaccard attention mechanism within SNNs to adapt to target domains without requiring source domain data. Similarly, He et al. (2024) facilitates the transfer of learned representations from static to dynamic event-based domains by adapting SNNs to process temporal information. Additionally, Zhan et al. (2024) converts RGB images into spike-based neuromorphic data, enabling SNNs to process visual information across various domains efficiently. However, the inherent difficulty of complex graph topologies makes it infeasible to apply spiking transfer learning to SGNs directly. To this end, we introduce a specialized domain adaptation method specifically tailored for SGNs.
@@ 3 Preliminaries
You et al. (2023) Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Redko et al. (2017), and the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, with probability at least $1-\delta$ the following inequality holds:
$\epsilon_{T}(h,\hat{h})\leq$ $\ \hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log\left(\frac{eN_{S}}{d}\right)+\frac{1}{N_{S}}\log\left(\frac{1}{\delta}\right)}$
$+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$. The first Wasserstein distance is defined as Villani et al. (2009): $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$.
@@
Spiking Graph Networks. In contrast to traditional artificial neural networks, SGNs Xu et al. (2021a); Zhu et al. (2022) convert input data into binary spikes over time, with each neuron in the SGNs maintaining a membrane potential that accumulates input spikes. A spike is produced as an output when the membrane potential exceeds a threshold, which is formulated as:
$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum_{j}w_{ij}\mathcal{A}\left(A,s_{\tau,j}\right)+b,\vspace{-4pt}$
$s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),$ (1)
where $\mathbb{H}(x)$ is the Heaviside function, which is the non-differentiable spiking function, $\mathcal{A}$ is the graph aggregation operation, and $A$ is the adjacency matrix of a graph. $s_{\tau,i}$ denotes the binary spike train of neuron $i$, and $\lambda$ is the constant. $w_{ij}$ and $b$ are the weights and bias of each neuron.
@@ 4.1 Degree-aware Personalized Spiking Representation
Specifically, we first $set$ all the degrees of nodes in the source domain graphs, i.e., $D^{s}=set(D_{1}^{s}\cup\cdots\cup D_{N_{s}}^{s})$, where $D_{i}$ denotes the degree set of graph $G_{i}^{s}$, and $set(\cdot)$ operation is an unordered sequence of non-repeating elements. Considering that low-degree nodes are more challenging to trigger while high-degree nodes trigger more easily, we propose setting higher thresholds for high-degree nodes and lower for low-degree nodes, which is formulated as:
$s_{\tau}^{d_{i}^{s}}=\mathbb{H}(u_{\tau}$ $-V_{th}^{d_{i}^{s}}),\quad S^{d_{i}^{s}}=avg(s_{\tau}^{d_{i}^{s}}),$ (2)
$V_{th}^{d_{i}^{s}}$ $=(1-\alpha)V_{th}^{d_{i}^{s}}+\alpha S^{d_{i}^{s}},$
where $V_{th}^{d_{i}^{s}}$ is the threshold of degree $d_{i}^{s}\in D^{s}$, initially set to $V_{th}$, and $\alpha$ is a hyper-parameter. The $avg(\cdot)$ operation takes the average of spiking representation with degree $d_{i}^{s}$. Consequently, high-degree nodes tend to achieve high $S^{d_{i}^{s}}$, which leads to an iterative increase in the threshold corresponding to degree $d_{i}^{s}$ and conversely for lower-degree nodes. To further explore the background mechanism of the hypothesis, we have the following analysis.
Hypothesis 1
In graph spiking networks, nodes with low-degree are more challenging to trigger while high-degree nodes trigger more easily.
The details analysis are introduced in Appendix A. With different thresholds for different node degrees, we can obtain the personalized node spiking representation $\mathbf{s}_{v\in G_{i}^{s}}^{d_{j}^{s}}$. Then, we summarize all node representations with a readout function into the graph-level representation and output the prediction with a multi-layer perception (MLP) classifier:
$\mathbf{s}_{i}=$ $\operatorname{READOUT}\left(\left\{\mathbf{s}_{v}^{d_{j}^{s}}\right\}_{v\in G_{i}^{s}}\right),$ (3)
$\hat{y}_{i}^{s}=H(\mathbf{s}_{i}),$
@@
However, in the scenario of domain adaptation, two significant issues remain unresolved. The first issue is that degrees in the target domain may be unseen in the source, rendering the thresholds ineffective for these degrees. The second is that the thresholds in the target may differ from those in the source, simply applying the source domain’s thresholds could lead to performance degradation. To alleviate the first issue, we initialize the threshold $V_{th}^{d_{i}^{t}}$ with the same value, where $d_{i}^{t}\notin D^{s}$. Then, with the training process of adversarial alignment, we iteratively update the threshold for degree $d_{i}^{t}$ with Eq. 2. To address the second issue, we incorporate the pseudo-label distillation module in Section 4.3 to guide the update of source degree thresholds on the target domain.